OpenStudy (anonymous):

Graph the linear equation. y = 3x – 4

OpenStudy (anonymous):

To graph a line like that, you first need to find the points at which the line intercepts the x and y axes.

OpenStudy (anonymous):

The point at which the line touches the y axis is called the y-intercept. This is the point where the value of x will be equal to zero. Given y = 3x -4 and x = 0, what is the value of y?

OpenStudy (anonymous):

Since you've got 2 points and a slope, all from the same line, use point-slope form

OpenStudy (anonymous):

Well.....to graph, all you need is two points. Easiest points to find are the x and y intercepts where the line touches the x and y axes respectively.

OpenStudy (anonymous):

Oh

OpenStudy (anonymous):

Its like drawing a line on paper. Mark two points, use a ruler to connect them. Same idea.

OpenStudy (anonymous):

The point at which the line touches the y axis is called the y-intercept. This is the point where the value of x will be equal to zero. Given y = 3x -4 and x = 0, what is the value of y?

OpenStudy (anonymous):

Can you tell me what the value of y would be in the above case?

OpenStudy (anonymous):

this is me i got blocked

OpenStudy (anonymous):

y = 3x - 4 Since x = 0, you have: y = 3(0) - 4 y = 0 - 4 y = -4 So, y-intercept is (0, -4). Similarly, for x intercept, you need to plug in y = 0. y = 3x - 4 0 = 3x - 4 3x = 4 x = 4/3 So, x-intercept is (4/3, 0). Now, you have two points on the graph (0, -4) and (4/3, 0). Connect them.

OpenStudy (anonymous):

whats the answer?

OpenStudy (anonymous):

You got to graph that line connecting those points.

OpenStudy (anonymous):

can you do it?

OpenStudy (anonymous):

hello

OpenStudy (anonymous):